use the pythagorean theorem to find the indicated quantity. Round your answer to the nearest hundredth.
10^2 = 6^2 + b^2

Answer by checkley77(12844)

10^2 = 6^2 + b^2
100=36+b^2
b^2=100-36
b^2=64
b=sqrt64
b=8 ans.